If a chord of circle x2 + y2 = 8 makes equal intercepts of length a on axes, then (1) |a| 4

Grade:11

1 Answers

Vikas TU
7 years ago
The chord of the circle let xx1 + yy1 – 8 =0
writing it in itnercept form we would get,
x/(y1/x1) + y/(-1) = -8/y1
and given intercepts are equal to a
hence, we get,
x1 = 2 and y1 =2
a = root(4 + 4) => 2root2 which is less than 4.